# DegDiT: Controllable Audio Generation with Dynamic Event Graph Guided Diffusion Transformer

### v1

@@ Abstract
Controllable text-to-audio generation aims to synthesize audio from textual descriptions while satisfying user-specified constraints, including event types, temporal sequences, and onset and offset timestamps. This enables precise control over both the content and temporal structure of the generated audio. Despite recent progress, existing methods still face inherent trade-offs among accurate temporal localization, open-vocabulary scalability, and practical efficiency. To address these challenges, we propose DegDiT, a novel dynamic event graph-guided diffusion transformer framework for open-vocabulary controllable audio generation. DegDiT encodes the events in the description as structured dynamic graphs. The nodes in each graph are designed to represent three aspects: semantic features, temporal attributes, and inter-event connections. A graph transformer is employed to integrate these nodes and produce contextualized event embeddings that serve as guidance for the diffusion model. To ensure high-quality and diverse training data, we introduce a quality-balanced data selection pipeline that combines hierarchical event annotation with multi-criteria quality scoring, resulting in a curated dataset with semantic diversity. Furthermore, we present consensus preference optimization, facilitating audio generation through consensus among multiple reward signals. Extensive experiments on AudioCondition, DESED, and AudioTime datasets demonstrate that DegDiT achieves state-of-the-art performances across a variety of objective and subjective evaluation metrics. Some generated examples are available at DegDiT Demo Page.
@@ I Introduction
Figure 1: Comparison between naive text-only guided (left) and our proposed dynamic event graph (DEG) guided (right) diffusion model generation. The naive text-only guided approach relies solely on textual input to drive the diffusion process. By integrating structured event graphs, our approach effectively captures temporal dependencies, enabling the generation of accurate and consistent timestamps.
@@
Building upon these advances, text-to-audio (TTA) generation [11, 12, 13] has emerged as a particularly promising application of diffusion models. TTA systems aim to synthesize realistic audio content from natural language descriptions, offering transformative potential for multimedia production, accessibility tools, and creative applications. Current state-of-the-art systems like AudioLDM [14, 15], Tango [12, 16] and TangoFlux [17] have demonstrated impressive capabilities in audio generation. Despite these advances, achieving fine-grained control over the generated audio remains a significant challenge for TTA systems. Such limitations hinder the use of TTA systems in applications that require detailed audio control. To solve this problem, one might consider drawing inspiration from controllable image generation methods [18, 19, 20], which have achieved remarkable success in manipulating visual content according to the user-specified conditions. However, these approaches cannot be directly adapted to audio due to modality differences, especially the temporal dynamics and continuous signal characteristics of audio data.
@@
Toward enabling precise control of sound events, recent works have explored distinct but complementary strategies. As the pioneer in conditional audio generation, MC-Diffusion [21] introduces multi-scale conditioning to simultaneously govern temporal and spectral attributes of generated audio. PicoAudio [22] addresses the precision of timestamps more aggressively by leveraging LLM to conduct a class-timestamp matrix as guidance. However, this approach is limited to a fixed set of classes due to the predefined dimensionality of the matrix. While AudioComposer [23] introduces a fully end-to-end natural language-based generation approach, it often suffers from missing audio events and degraded output quality due to its over-reliance on text descriptions. Most recently, FreeAudio [24] introduces a training-free method using aggregation of preliminary and timestamp-independent attention maps, which introduces operational complexity due to its reliance on carefully designed parameters.
@@
To address these issues, we propose DegDiT, which models sound events within the multi-event audio prompt as temporally dynamic graphs that evolve across the audio timeline, as depicted in Fig. 1. Each graph comprises four node categories that collectively represent the event’s spatio-temporal signature: (1) event category, (2) frame-level occurrence, (3) temporal boundary attributes (localized onset/offset), and (4) inter-event relations. A graph transformer processes these dynamic graphs, aggregating the spatio-temporal context through attention mechanisms to produce contextualized embeddings. These enriched embeddings are then fused with the textual features and provided as guidance for the diffusion model.
@@
Considering the multi-dimensional evaluation requirements in controllable audio generation, we propose Consensus Preference Optimization (CoPO), a reinforcement learning framework that integrates diverse reward signals for comprehensive preference modeling. Unlike conventional binary preference learning, CoPO leverages quantitative preference intensities derived from comprehensive quality assessments, including text alignment, event alignment, temporal accuracy, and audio quality. By aggregating these reward components with learnable weights, CoPO captures the consensus among multiple reward signals, facilitating more nuanced model optimization.
@@ II-B Controllable Image Generation
To achieve precise spatial control in the image generation era, researchers employ auxiliary control signals to guide the diffusion process. ControlNet [18] utilizes structural cues such as depth maps to constrain the generation process, ensuring output images adhere to predefined spatial configurations. Subsequent works, including IP-Adapter [30], PulID [19], and ConsisID [20], have further advanced this approach by incorporating dedicated image encoders to extract spatial features from the reference images. Beyond these trained adapters, training-free methods have emerged to enhance flexibility and reduce computational overhead. Regional Prompting [31] achieves spatial control using region masks in pre-trained models, and RPG [32] leverages multi-modal LLMs to decompose complex prompts for compositional generation.
@@ III Method
In this section, we introduce DegDiT, a novel framework for fine-grained, controllable text-to-audio generation based on dynamic event graphs. The goal of DegDiT is to synthesize audio that not only matches the semantic content of the text prompt but also adheres to detailed temporal and structural specifications of sound events. As shown in Fig. 2, the overall framework constructs dynamic event graphs to represent audio events along with their temporal characteristics and relationships. These representations are then processed by a transformer to capture both contextual information. Finally, a multi-modal diffusion transformer generates the output audio conditioning on both the text prompt and the structured graphs, ensuring the generated content maintains semantic and temporal coherence with the original specifications.
@@ III-A Dynamic Event Graph Representation
Given a text prompt $T$ and a set of event descriptions $E$, DegDiT constructs a collection of dynamic event graphs $\mathcal{G}=\{\mathcal{G}_{1},\mathcal{G}_{2},\ldots,\mathcal{G}_{N}\}$, where each graph $\mathcal{G}_{i}=(\mathcal{V}_{i})$ represents a unique audio event with its characteristics as nodes. The node set $\mathcal{V}_{i}$ consists of three components that capture different aspects of the event: an event node $\mathbf{E}_{i}$ encoding the basic attributes, a temporal activation node $\mathbf{F}_{i}$ tracking frame-level presence, and a relational node $\mathbf{A}_{i}$ modeling interactions with other event nodes.
@@
The event node $\mathbf{E}_{i}$ serves as the anchor point, capturing the core attributes of each audio event through the tuple $(c_{i},s_{i},e_{i},\alpha_{i})$, where $c_{i}$ identifies the event category (such as “dog barking”), while $s_{i}$ and $e_{i}$ mark its start and end times in seconds, and $\alpha_{i}\in[0,1]$ represents its intensity. This compact representation forms the basis for both semantic understanding and temporal localization of the events. The temporal activation node $\mathbf{F}_{i}$ tracks how the event manifests across time, which we formalize through a feature matrix $\mathbf{F}\in\mathbb{R}^{N\times F}$ that provides fine-grained control over event timing. Here, each element $\mathbf{F}_{i,j}$ indicates how prominently event $i$ appears in time frame $j$, calculated by the proportion of duration within the frames, precisely aligning the generated audio and the specified frame. The relational node $\mathbf{A}_{i}$ encodes how events interact temporally, supporting relationships like before, after, overlaps, contains, and contained-by. These interactions are systematically captured in an adjacency tensor $\mathbf{A}\in\mathbb{R}^{N\times N\times 5}$, where each slice $\mathbf{A}_{i,j,:}$ measures the strength of temporal relationship between events $i$ and $j$, such as their degree of overlap or containment, allowing the model to reason about complex event co-occurrences.
@@
To encode the textual description of the events, traditional approaches [22] often rely on a fixed set of event types, limiting their ability to generalize to novel or rare events. In contrast, we employ an open-vocabulary text encoder based on a pretrained FLAN-T5 large language model [33], allowing flexible encoding of arbitrary event descriptions while leveraging the model’s rich semantic knowledge. For each event, the type description $c_{i}$ is first tokenized and passed through the FLAN-T5 encoder:
$\mathbf{h}^{\text{type}}_{i}=\textrm{T5Encoder}(c_{i}),$ (1)
@@
where $\mathbf{h}^{\text{type}}_{i}$ is a sequence of token embeddings that capture the nuanced meaning of the event description. To obtain a fixed-length representation suitable for graph processing, we apply mean pooling over the token dimension, followed by a linear projection that adapts the embedding space to our specific task:
$\mathbf{e}^{\text{type}}_{i}=\textrm{Proj}(\textrm{MeanPool}(\mathbf{h}^{\text{type}}_{i})).$ (2)
@@
This approach enables the model to capture the semantic information of the event types while maintaining the flexibility to handle unseen event categories during inference.
@@
After acquiring the semantic representation of each event, precise temporal modeling addresses the temporal characteristics of the sound event, specified by its onset and offset timestamps $(s_{i},e_{i})$, which are processed through a dedicated encoding pathway. We implement this via a multi-layer perceptron (MLP) with sinusoidal activation functions that are particularly suited for representing periodic temporal patterns:
$\mathbf{t}_{i}=\textrm{MLP}([s_{i},e_{i}]).$ (3)
@@
The final node representation $\mathbf{g}^{(0)}_{i}$ is obtained through element-wise addition of the semantic and temporal embeddings, combining both aspects of event understanding:
$\mathbf{g}^{(0)}_{i}=\mathbf{e}^{\text{type}}_{i}+\mathbf{t}_{i}.$ (4)
The resulting node embeddings capture both the semantic meaning and temporal occurrence of each sound event.
Modeling the dependencies and interactions among multiple events is essential for generating coherent and contextually appropriate audio that reflects real-world sound relationships. Each event node $\mathcal{G}_{i}$ aggregates contextual information from other events using the relation tensor:
$\mathbf{r}_{i}=\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}\textrm{RelationEncoder}(\mathbf{A}_{i,j,:}),$ (5)
where $\mathcal{N}_{i}$ denotes the set of all other events, and $\mathbf{A}_{i,j,:}$ encodes the multi-relational edge features between $v_{i}$ and $v_{j}$, allowing the model to learn different weights for different types of temporal relationships. This aggregation allows the model to reason about both the local context and global temporal dependencies of each event.
@@
In parallel, the temporal activity pattern for the pre-defined frames and each event is encoded to capture the detailed evolution of sound events over time:
@@
$\mathbf{f}_{i}=\alpha_{i}\cdot\textrm{FrameEncoder}(\mathbf{F}_{i,:}),$ (6)
@@
where $\mathbf{F}_{i,:}$ is the temporal activation vector for event $i$ across all pre-defined frames, and the intensity weighting $\alpha_{i}$ preserves the relative importance of different events. We utilize these encoding features to maintain precise alignment between the generated audio and the specified temporal structure throughout the generation process.
Finally, a learned transformation integrates the node features with relational and temporal embeddings to produce the union feature representation:
$\mathbf{g}^{(1)}_{i}=\textrm{Fusion}([\mathbf{g}^{(0)}_{i},\mathbf{r}_{i}+\mathbf{f}_{i}]).$ (7)
@@
These representations are then processed by a transformer encoder, which models both local and global dependencies among events through its self-attention mechanism:
@@
where $\mathbf{G}^{(1)}$ stacks all event node features, PosEmb provides positional information about the event sequence, and $\mathbf{M}$ is an attention mask indicating valid events. The output $\mathbf{H}$ represents a set of dynamic event graph embeddings that comprehensively capture the structure and dynamics of the entire audio scene, with dimensionality matching the text embeddings in the diffusion model for seamless integration.

### v2

@@ Abstract
Controllable text-to-audio generation aims to synthesize audio from textual descriptions while satisfying user-specified constraints, including event types, onset and offset timestamps, and temporal sequences. This enables precise control over both the content and temporal structure of the generated audio. Despite recent progress, existing methods still face inherent trade-offs among accurate temporal localization, open-vocabulary scalability, and practical efficiency. To address these challenges, we propose DegDiT, a novel dynamic event graph-guided diffusion transformer framework for open-vocabulary controllable audio generation. DegDiT encodes the events from the text description into structured dynamic graphs. The nodes represent distinct audio events, while edges encode the temporal relationships between them. A transformer is employed to integrate these graphs and produce contextualized event embeddings that serve as guidance for the diffusion model. To ensure high-quality and diverse training data, we introduce a quality-balanced data selection pipeline that combines hierarchical event annotation with multi-criteria quality scoring, resulting in a curated dataset with semantic diversity. Furthermore, we present consensus preference optimization, facilitating audio generation through consensus among multiple reward signals. Extensive experiments on AudioCondition, DESED, and AudioTime datasets demonstrate that DegDiT achieves state-of-the-art performances across a variety of objective and subjective evaluation metrics. Some generated examples are available at DegDiT Demo Page.
@@ I Introduction
Figure 1: Comparison between naive text-only guided (left) and our proposed dynamic event graph guided (right) diffusion model generation. The naive text-only guided approach relies solely on textual input to drive the diffusion process. By integrating structured event graphs, our approach effectively captures temporal dependencies, enabling the generation of accurate and consistent timestamps.
@@
Building upon these advances, text-to-audio (TTA) generation [22, 16, 54] has emerged as a particularly promising application of diffusion models. TTA systems aim to synthesize realistic audio content from natural language descriptions, offering transformative potential for multimedia production, accessibility tools, and creative applications. Current state-of-the-art systems like AudioLDM [33, 34], Tango [16, 36] and TangoFlux [23] have demonstrated impressive capabilities in audio generation. Despite these advances, achieving fine-grained control over the generated audio remains a significant challenge for TTA systems. Such limitations hinder the use of TTA systems in applications that require detailed audio control. To solve this problem, one might consider drawing inspiration from controllable image generation methods [58, 19, 57], which have achieved remarkable success in manipulating visual content according to the user-specified conditions. However, these approaches cannot be directly adapted to audio due to modality differences, especially the temporal dynamics and continuous nature of audio signals.
@@
Toward enabling precise control of text-to-audio generation, recent works have explored various strategies. As an early exploration, MC-Diffusion [18] introduces multi-scale conditioning to simultaneously govern temporal and spectral attributes of generated audio. PicoAudio [51] addresses the precision of timestamps by leveraging LLM to conduct a class-timestamp matrix as guidance. However, this approach is limited to a fixed set of classes due to the predefined dimensionality of the matrix. While AudioComposer [47] introduces a fully end-to-end natural language-based generation approach, it often suffers from missing audio events and degraded output quality due to its over-reliance on text descriptions. Most recently, FreeAudio [25] introduces a training-free method using aggregation of preliminary and timestamp-independent attention maps, which introduces operational complexity due to its reliance on carefully designed parameters.
@@
To address these issues, we propose DegDiT, which models sound events within the multi-event audio prompt as temporally dynamic graphs that evolve across the audio timeline, as depicted in Fig. 1. In the graphs, each event is represented as a node carrying categories, temporal activation, and boundary attributes, while the edges between nodes encode their temporal relationships. A graph transformer processes these dynamic graphs, aggregating the spatio-temporal context through attention mechanisms to produce contextualized embeddings. These enriched embeddings are then fused with the textual features and provided as guidance for the diffusion model.
@@
Considering the multi-dimensional evaluation requirements in controllable audio generation, we propose Consensus Preference Optimization (CoPO), a reinforcement learning framework that integrates diverse reward signals for preference modeling. Unlike conventional binary preference learning, CoPO leverages quantitative preference intensities derived from quality assessments across multiple dimensions: text alignment, event alignment, temporal accuracy, and audio quality. By aggregating these reward components with learnable weights, CoPO captures the consensus among multiple reward signals, facilitating more nuanced model optimization.
@@ II-B Controllable Image Generation
To achieve precise spatial control in the image generation era, researchers employ auxiliary control signals to guide the diffusion process. ControlNet [58] utilizes structural cues such as depth maps to constrain the generation process, ensuring output images adhere to predefined spatial configurations. Subsequent works, including IP-Adapter [56], PulID [19], and ConsisID [57], have further advanced this approach by incorporating dedicated image encoders to extract spatial features from the reference images. Beyond these trained adapters, training-free methods have emerged to enhance flexibility and reduce computational overhead. Regional Prompting [5] achieves spatial control using region masks in pre-trained models, and RPG [55] leverages multi-modal large language models (MLLMs) to decompose complex prompts.
@@ II-C Controllable Audio Generation
II-D Graph-based Representation Learning
Graph neural networks [41, 4, 49] provide powerful frameworks for structured data modeling. [29] introduced graph convolutional networks for neighbor aggregation, while [45] developed graph attention networks with adaptive weighting mechanisms. Beyond GNNs, [28] demonstrated that pure transformer models can learn graph representations when equipped with structural encodings. In visual scene understanding, graph representations model complex relationships effectively. [53] proposed iterative message passing for scene graph generation, and [24] introduced Action Genome, representing actions as spatio-temporal scene graphs for dynamic scene understanding. Capitalizing on its inherent strengths in modeling entities and their relational structure, we leverage graph-based representation learning for controllable audio generation in this paper.
@@ III Method
In this section, we introduce DegDiT, a novel framework for fine-grained, controllable text-to-audio generation based on the dynamic event graphs. The goal of DegDiT is to synthesize audio that not only matches the semantic content of the text prompts but also adheres to detailed temporal specifications of sound events. As shown in Fig. 2, the overall framework constructs dynamic event graphs to represent audio events along with their temporal characteristics and relationships. These representations are then processed by a transformer to capture the contextual information. Finally, we train a multi-modal diffusion transformer that generates audio conditioned on both text prompts and graph representations, ensuring semantic and temporal alignment with the input specifications.
@@ III-A Dynamic Event Graph Representation
Given a text prompt $T$, DegDiT constructs a dynamic event graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where:
  • $\mathcal{V}=\{v_{1},v_{2},\ldots,v_{N}\}$ is the set of all nodes representing audio events,
  • $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges representing interactions between events.
@@
The node set $\mathcal{V}$ consists of $N$ audio events, where each $v_{i}\in\mathcal{V}$ is associated with a pair of specialized properties: an event property $\mathbf{E}_{i}$ and a temporal activation property $\mathbf{F}_{i}$. The event property $\mathbf{E}_{i}$ serves as the anchor point, capturing the core attributes of each audio event through the tuple $(c_{i},s_{i},e_{i},\alpha_{i})$, where $c_{i}$ identifies the event category (such as “dog barking”), while $s_{i}$ and $e_{i}$ mark its start and end times in seconds, and $\alpha_{i}\in[0,1]$ represents its intensity. This compact representation forms the basis for both semantic understanding and temporal localization of the events. On the other hand, the temporal activation property $\mathbf{F}$ tracks how the event manifests across time, which we formalize through a feature matrix $\mathbf{F}\in\mathbb{R}^{N\times F}$ that provides fine-grained control over event timing. Here, each element $\mathbf{F}_{i,j}$ indicates how prominently event $i$ appears in time frame $j$, calculated by the proportion of duration within the frames. Additionally, we define edges in $\mathcal{E}$ to capture temporal interactions between events, supporting five relationship types: before, after, overlaps, contains, and contained-by. These interactions are systematically captured in an adjacency tensor $\mathbf{A}\in\mathbb{R}^{N\times N\times 5}$, where each slice $\mathbf{A}_{i,j,:}$ measures the relationship between events $i$ and $j$.
@@
To encode the textual description of the events, traditional approaches [51] often rely on a fixed set of event types, limiting their ability to generalize to novel or rare events. In contrast, we employ an open-vocabulary text encoder based on a pretrained FLAN-T5 large language model [8], allowing flexible encoding of arbitrary textual event descriptions while leveraging the model’s rich semantic knowledge. For each event, the category $c_{i}$ is firstly tokenized and passed through the text encoder:
$\mathbf{h}^{\text{cat}}_{i}=\textrm{T5Encoder}(c_{i})\in\mathbb{R}^{L\times d_{\text{T5}}},$ (1)
@@
where $\mathbf{h}^{\text{cat}}_{i}$ is a sequence of token embeddings that capture the nuanced meaning of the event category, with $L$ being the sequence length and $d_{\text{T5}}$ the hidden dimension of the text encoder. To obtain a fixed-length representation suitable for graph processing, we apply mean pooling along the sequence dimension, followed by a linear projection that adapts the embedding space to our specific task:
$\mathbf{e}^{\text{cat}}_{i}=\textrm{Proj}(\textrm{MeanPool}(\mathbf{h}^{\text{cat}}_{i}))\in\mathbb{R}^{d_{\text{emb}}}.$ (2)
@@
This enables the model to capture the semantic information of the event categories while maintaining the flexibility to handle unseen event categories during inference.
@@
After acquiring the semantic representation of each event, precise temporal modeling addresses the temporal characteristics of the sound event, specified by its onset and offset timestamps $(s_{i},e_{i})$. We implement this via a multi-layer perceptron (MLP) with sinusoidal activation functions that are particularly suited for representing periodic temporal patterns:
$\mathbf{t}_{i}=\textrm{MLP}(\text{Concat}(s_{i},e_{i}))\in{\mathbb{R}^{d_{\text{emb}}}}.$ (3)
@@
The final event representation $\mathbf{g}^{(0)}_{i}$ is obtained through element-wise addition of the semantic and temporal embeddings, combining all the aspects of event property attributes:
$\mathbf{g}^{(0)}_{i}=\mathbf{e}^{\text{cat}}_{i}+\mathbf{t}_{i}\in\mathbb{R}^{d_{\text{emb}}}.$ (4)
@@
In parallel, the temporal activity property for the pre-defined frames and each event is also encoded to capture the fine-grained temporal structure of audio events:
@@
$\mathbf{f}_{i}=\alpha_{i}\cdot\textrm{FrameEncoder}(\mathbf{F}_{i,:})\in\mathbb{R}^{d_{\text{frame}}},$ (5)
@@
where $\mathbf{F}_{i,:}\in\mathbb{R}^{F}$ is the temporal activation vector for event $i$ across all $F$ pre-defined frames, $\textrm{FrameEncoder}:\mathbb{R}^{F}\to\mathbb{R}^{d_{\text{frame}}}$ projects the frame-level features, and the intensity weighting $\alpha_{i}$ preserves the relative importance of different events. These features maintain precise alignment between the specified temporal structure and the generated audio.
Another key point in controllable audio generation lies in capturing the interplay between multiple sound events, which serves as the mechanism to achieve the coherence of audio events. To capture these inter-event dependencies, we leverage the edge set $\mathcal{E}$ and its representation in the adjacency tensor $\mathbf{A}$. Each event node $v_{i}$ aggregates contextual information from others through:
$\mathbf{r}_{i}=\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}\textrm{RelationEncoder}(\mathbf{A}_{i,j,:})\in\mathbb{R}^{d_{\text{rel}}},$ (6)
where $\mathcal{N}_{i}$ denotes the set of all other events, $\mathbf{A}_{i,j,:}\in\mathbb{R}^{5}$ encodes the edge features between $v_{i}$ and $v_{j}$, and $\textrm{RelationEncoder}:\mathbb{R}^{5}\to\mathbb{R}^{d_{\text{rel}}}$ is a learnable mapping. We set $d_{\text{rel}}=d_{\text{frame}}$ to ensure dimensional compatibility for subsequent fusion operations.
Finally, a learnable lightweight feed-forward network $\textrm{Fusion}(\cdot)$ integrates the event property features with relation and frame embeddings to produce the union representation:
$\mathbf{g}^{(1)}_{i}=\textrm{Fusion}\big(\text{Concat}(\mathbf{g}^{(0)}_{i},\mathbf{r}_{i}+\mathbf{f}_{i})\big)\in{\mathbb{R}^{d_{\text{g}}}}.$ (7)
@@
The collected representation of all the events $\mathbf{G}^{(1)}=[\mathbf{g}^{(1)}_{1};\mathbf{g}^{(1)}_{2};\dots;\mathbf{g}^{(1)}_{N}]\in\mathbb{R}^{N\times d_{g}}$ is then processed by a transformer encoder, modeling both local and global dependencies among events through the self-attention mechanism:
@@
where $\textrm{PosEmb}\in\mathbb{R}^{N\times d}$ adds positional information, and $\mathbf{M}$ is an optional attention mask. The output $\mathbf{H}$ is a set of dynamic event graph embeddings that encode the structure and dynamics of the entire audio scene. Crucially, its embedding dimension aligns with the text embedding dimension of the diffusion model.